Nina can wash the walls in a certain room in 2 hours and mark can wash these walls in 5 hours. How long would it take them to complete the task if they work together?
Answer by addingup(3677)

In 1 hour Nina can wash 1/2
In 1 hour Mark can wash 1/5
Find common denominator:
1/2 multiply times 5 = 5/10
1/5 multiply times 2 = 2/10
5/10+2/10 = 7/10
1 hour has 60 minutes:
7/10*60/1 = 420/10 = 42 minutes is your answer